# CDR: Customizable Density Ratios of Strong-over-weak LLMs for Preference Annotation

### v2

@@ Abstract
Preference tuning of large language models (LLMs) relies on high-quality human preference data, which is often expensive and time-consuming to gather. While existing methods can use trained reward models or proprietary model as judges for preference annotation, they have notable drawbacks: training reward models remain dependent on initial human data, and using proprietary model imposes license restrictions that inhibits commercial usage. In this paper, we introduce customized density ratio (CDR), a training-free and highly effective method that leverages off-the-shelf LLMs for preference data annotation. Our approach uses the log-density ratio between a better-aligned LLM and a less aligned LLM as a reward signal. We explores 221 different LLMs pairs and empirically demonstrate that increasing the performance gap between paired LLMs correlates with better reward generalization. Furthermore, we show that tailoring the density ratio reward function with specific criteria and preference exemplars enhances performance across domains and within target areas.
In our experiment using density ratio from a pair of Mistral-7B models, CDR achieves a RewardBench score of 82.6, outperforming the best trained reward functions from same model class and demonstrating competitive performance against SoTA models in Safety (91.0) and Reasoning (88.0) domains. We use CDR to annotate an on-policy preference dataset with which we preference tune Llama-3-8B-Instruct with SimPO. Using reward signals from two relatively weak models, our approach pushes Llama-3-8B to achieve a 37.4% ($+$15.1%) win rate on ArenaHard and a 40.7% ($+$17.8%) win rate on Length-Controlled AlpacaEval 2.0, along with a score of 8.0 on MT-Bench.
Figure 1: Scatter plot of model alignment gap (x-axis) versus RewardBench score (y-axis) for models in the Llama-8B (orange) and Mistral-7B (blue) families. Models’ human-alignment gap is measured using difference in ArenaHard scores between $\pi_{\text{strong}}$ and $\pi_{\text{weak}}$ policies belonging to the same model family. Each point represents one of 221 unique model pairings from these families (10 models from Llama-8B, 11 from Mistral-7B). Models include Base, SFT, and preference fine-tuned with algorithms such as DPO, PPO, KTO, RRHF, ORPO, SimPO, IPO, and SLiC-HF. The plot demonstrates a positive correlation between model alignment gap and RewardBench score, indicating that higher alignment gaps correspond with more effective reward model performance.
@@ 1 Introduction
Preference tuning has advanced the capabilities of large language models (LLMs), but this progress relies on high-quality human preference data which is both costly and time-consuming to gather. Cutting-edge models (e.g., ChatGPT, GPT-4, Claude-3) are aligned with curated, quality-controlled human preference data, typically provided by specialized companies. While effective, this approach limits broader adoption due to prohibitive costs and limited transparency in data collection (Wang et al., 2024c). AI-feedback solutions are emerging as an alternative—either through a trained reward model (Dong et al., 2024) or proprietary LLM-as-a-judge (Cui et al., 2023). However, training reward models still rely on costly initial human preference data, and proprietary LLM-as-a-judge approaches introduce licensing restrictions that generally prevent commercial use.
This paper introduces customized density ratio (CDR) that leverages the density ratio between off-the-shelf open-source LLMs to efficiently facilitate preference data annotation. Our method uses the log-density ratio between a better-aligned model and less-aligned model to annotate preference data. We show that a higher alignment gap between model pairs yields improved preference signal. This observation, referred to as the “Strong-over-Weak Hypothesis”, is supported by our experiments across 221 model combinations (Figure 1). Notably, log-density ratios between a post-DPO model and a pre-DPO model, known as the DPO implicit reward (Rafailov et al., 2023), have not gained widespread adoption due to limitations in generalizability and high reward variance in reward performance across different model choices (Lambert et al., 2024; Lin et al., 2024). We show that the implicit DPO reward is an empirically suboptimal special case of the strong-over-weak density ratio reward, with its performance variance reducible by consistently choosing a weaker model as the reference (Figure 2). Our findings highlight the importance of selecting model pairs with a sufficient alignment gap and demonstrate the flexibility to construct the density ratio reward using models trained with various objectives (such as SFT, RRHF, SLiC-HF, ORPO, SimPO, KTO, IPO, etc.)
We customize our density-ratio reward function to align with the domain of each sample in the annotation set. Given that human preferences span multiple dimensions (e.g., trustworthiness, reliability), an effective reward function should adapt to requirements specific to each domain. Our CDR introduces an end-to-end process, from identifying the domain of each user query to customizing the reward function to prioritize the relevant preference criteria. Specifically, CDR first uses an adaptive router to identify the domain of each user query (e.g., chat, reasoning, safety). It then applies domain-specific instructions and in-context learning (ICL) examples to clarify preference criteria. In this way, we customize a density-ratio reward function from a general preference signal to domain-specific annotators. Experimental results demonstrate that adaptively customized density ratio significantly improve in both overall and target domain reward generalization.
The main contributions of this paper are as follows.
  • Choosing Models via Strong-over-Weak Hypothesis. We propose a training-free framework to leverage the density ratio between a better-aligned LLM and a less-aligned LLM as a reward signal for annotating preference data. We introduce the “Strong-over-Weak Hypothesis”, which suggests increasing the preference gap between the two LLMs to improve the accuracy of the density ratio reward function. Through extensive experimentation on 221 model pairs, we empirically validate this hypothesis. Our findings can reduce the reward variance seen in existing density ratio methods (e.g., DPO Implicit reward) and offer guidance on selecting effective model pairs for density ratio-based reward functions.
  • Customizing Reward Function via Prompting. We customize the density ratio reward function for target domains by incorporating domain-specific instructions and ICL examples. Our experiment on RewardBench shows significant domain-wise improvement after applying customized conditioning: Safety domain improved from 82.4 to 91.0, the Reasoning domain performance from 73.8 to 88.0, and the ChatHard domain from 60.4 to 69.7. CDR uses a LLM-based router to assign customized instructions for different examples in the annotation set. The adaptive CDR improves density ratio reward w/o prompting by 5.3 points overall on RewardBench, and exceeds best trained classifier of same model class.
  • Alignment Improvement. We use CDR of a pair of Mistral-7B models to annotate on-policy preference data collected through Best-of-N sampling. Training on this data, the Llama-3-8B-Instruct model is aligned to achieve a 37.4% (+15.1%) win rate on ArenaHard, a 40.7% (+17.8%) length-controlled win rate on AlpacaEval 2.0, and a score of 8.0 on MT-Bench, reaching performance comparable to SoTA trained reward functions.
@@ 2 Background
Before diving into preference data annotation, we review two core optimization objectives in preference tuning. The first optimization is a maximum likelihood estimator of the Bradley-Terry model that aims to learn an optimal reward function. Given dataset of $N$ samples $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, the objective in equation 1 aims to increase the reward for the preferred response $y_{w}^{(i)}$ while reducing it for the dispreferred response $y_{l}^{(i)}$ for each prompt $x^{(i)}$. The second optimization, equation 2, given a reward function $r$, focuses on finding an optimal policy $\pi$, e.g. LLM, that maximizes the reward while remaining close to a reference model $\pi_{\text{ref}}$.
$\operatorname*{arg\,min}_{r}-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(r(x,y_{w})-r(x,y_{l})\right)\right]\quad\quad\quad\quad$ (optimal reward) (1)
$\operatorname*{arg\,max}_{\pi}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi(y|x)}\left[r(x,y)\right]-\beta D_{\mathrm{KL}}\left[\pi(y|x)\|\pi_{\text{ref}}(y|x)\right]$ (optimal policy) (2)
where $\sigma$ is the sigmoid function. For preference tuning through RLHF (Ouyang et al., 2022), the process begins by optimizing (1) to identify an optimal reward function. This learned reward function is then incorporated into the second optimization step (2) to align the language model’s policy. Policy optimization (2) is typically solved using RL algorithms, such as PPO, because external reward is non-differentiable w.r.t model parameter.
Now, given a reward function $r(x,y)$, Rafailov et al. (2023) shows that the solution to equation 2 is
$\pi_{r}^{*}(y|x)=\frac{1}{Z(x)}\pi_{\text{ref}}(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right),$ (3)
which implies the corresponding reward function can be written in forms of the policies as:
$r(x,y)=\beta\log\left(\frac{\pi_{r}^{*}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}\right)+\beta\log Z(x).$ (4)
A key insight of DPO (Rafailov et al., 2023) is that this implicit reward function can then be incorporated into the reward optimization objective (1) to formulate a maximum likelihood objective for a parametrized policy $\pi_{\theta}$ directly, without explicitly learning a reward function. The DPO loss is as below
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\theta_{\text{ref}}}(y_{l}\mid x)}\right)\right)\right].$ (5)
DPO Implicit Reward
In this way, DPO objective in equation 5 skips the need of explicit reward and directly optimizes the parametric model $\pi_{\theta}$, which is equivalent to fitting a reparametrized BradleyTerry reward model in equation 1 under a change of variables (see Section 5 in Rafailov et al. (2023) for details). In other words, optimizing DPO objective also learns an implicit reward function $r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$ that approximates the ground-truth reward according to equation 1. Based on this connection, follow-up works (Lambert et al., 2024; Lin et al., 2024; Chen et al., 2024) directly leverage the implicit reward function to annotate preference data. Their data annotation process works as follow. For a given prompt $x$ and two responses $y_{1}$, $y_{2}$, the preference is annotated by comparing $r(x,y_{1})$ with $r(x,y_{2})$: the one with a higher reward is labeled as preferred, while the other is marked as dispreferred. This comparison can be simplified to compare the log-density ratios only, essentially
$\log\frac{\pi_{\theta}(y_{1}|x)}{\pi_{\text{ref}}(y_{1}|x)}\quad\text{v.s.}\quad\log\frac{\pi_{\theta}(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x)}.$
@@ 3 Method
We study two research questions critical to reward function design based on density ratio. First, we discuss how to find successful pairs of LLMs to construct effective density ratio reward (section 3.1). Our results reveal a strong correlation between the alignment gap of model pairs (measured by the ArenaHard score) and the effectiveness of the reward function (evaluated through the RewardBench score). These results also suggest that a wider range of model pairs beyond DPO implicit reward to be used for constructing reward functions, provided the gap in their human-aligned level is sufficient. Second, we study how to make density ratio a customizable and refined preference signal tailored to defined criterion (section 3.2). We show that using domain-specific instructions and exemplars to condition density ratio significantly improve overall and domain reward performance, without needing additional training.
@@
3.1 Reward Function Design using Log-Density Ratio
@@
We explore constructing density ratio reward functino with various pairings of LLMs. At first glance, one might assume that the DPO model and its reference policy would be the optimal pair for this purpose. To examine this hypothesis, we conduct an experiment applying online iterative DPO (Xiong et al., 2023; Xu et al., 2023; Swamy et al., 2024) to both the Mistral and Llama-3 model families. The key ideas of online iterative DPO training are: (1) the reference model is updated at each iteration (i.e., $\pi_{\text{ref}}=\pi_{\theta_{t-1}}$), and (2) the training data is also updated iteratively, with $y_{w}$ and $y_{l}$ generated by sampling from $\pi_{\theta_{t-1}}(\cdot\mid x)$ and annotated with preferences using an external reward function. We revisit the loss function in comparison to the DPO loss function in (5):
$\mathcal{L}_{\text{iter\_DPO}}(\pi_{\theta_{t}};\pi_{\theta_{t-1}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{t}}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta_{t}}(y_{w}\mid x)}{\pi_{\theta_{t-1}}(y_{w}\mid x)}-\log\frac{\pi_{\theta_{t}}(y_{l}\mid x)}{\pi_{\theta_{t-1}}(y_{l}\mid x)}\right)\right)\right].$ (6)
We ensure that the DPO objective is effectively optimized at each iteration, as we see consistent improvement in updated model’s ArenaHard score (see Figure 2). This process starts with an SFT model trained from Base, and three DPO checkpoints from 3 DPO optimization iterations (obtained from  Chen et al. (2024)). We use the log-density ratio between all model combinations within each family (Mistral or Llama-3) to define reward functions and evaluate the effectiveness of these reward functions via RewardBench.
In this online iterative DPO setting, the DPO model ($\pi_{\theta_{t}}$) at iteration $t$ uses the policy model at the previous iteration ($\pi_{\theta_{t-1}}$) as its reference model. According to the implicit DPO reward theory, one might expect the density ratio between $\pi_{\theta_{t}}$ and $\pi_{\theta_{t-1}}$ to provide an optimal reward function. However, Figure 2 shows that using weaker models–such as the base or SFT models—as the denominator in the log-density ratio, rather than $\pi_{\theta_{t-1}}$, leads to significantly better reward functions evaluated by RewardBench. This result shows that empirically, DPO implicit reward is suboptimal compared with simply choosing weaker reference policies, implication of which motivates us to propose “Strong-over-Weak Hypothesis“ to guide model pairing.
@@
$r(x,y)=\log\frac{\pi_{\text{strong}}(y\mid x)}{\pi_{\text{weak}}(y\mid x)}=\log\pi_{\text{strong}}(y\mid x)-\log\pi_{\text{weak}}(y\mid x).$ (7)
@@
In Section 4.1, we conduct extensive experiments on density ratio reward involving $221$ distinct model pairs. Our findings reveal a strong correlation between the alignment gap of $\pi_{\text{strong}}$ and $\pi_{\text{weak}}$ and the effectiveness of the reward function, as quantified by the RewardBench score. As shown in Figure 1, achieving an effective reward function in Eq. (7) necessitates a substantial human-alignment level difference between $\pi_{\text{strong}}$ and $\pi_{\text{weak}}$. We term this insight the “Strong-over-Weak Hypothesis" , which can serve as a guiding principle for selecting models in density ratio reward computation. Furthermore, our experiments span a range of models fine-tuned using different preference tuning strategies (e.g., DPO, SimPO, KTO, ORPO) to show that density ratio reward is not exclusive for DPO models, with further details provided in Figure 5. We summarize our key observations below.
@@
You are a helpful AI assistant. You follow the following guidelines when answering user questions. 1. Answer Constructive, Clear Questions - Provide an answer when the user asks for factual information, constructive advice, or help with personal growth. Focus on offering practical, positive guidance. 2. Recognize Jokes, Puns, and Fictional Contexts - Respond playfully when the question references humor, games, movies, or fictional scenarios. Acknowledge the fictional nature while keeping the tone light. 3. Avoid Answering Harmful, Illegal, or Malicious Questions - Do not engage if the question promotes harm, illegal activities, or unethical behavior. Politely but firmly refuse to provide an answer, while keeping the response respectful. 4. Handle Sensitive Topics with Empathy - Respond with care to questions about mental health, personal relationships, or emotionally charged situations. Acknowledge the user’s feelings, and offer general advice or suggest professional resources.

### v3

# Dr. SoW: Density Ratio of Strong-over-weak LLMs for Reducing the Cost of Human Annotation in Preference Tuning

@@ Abstract
Preference tuning relies on high-quality human preference data, which is often expensive and time-consuming to gather. In this paper, we introduce Dr.SoW (Density Ratio of Strong over Weak) a cost-effective method that eliminates the reliance for human annotation by leveraging off-the-shelf LLMs for preference data annotation. Dr.SoW uses the log-density ratio between a better-aligned and a less-aligned LLM as a reward signal. We evaluate Dr.SoW across 221 different LLM pairs and empirically find a strong correlation between the performance gap of the paired models and the quality of the reward signal. This insight provides a practical guideline for selecting LLMs for data annotation. Additionally, we introduce an end-to-end pipeline that customizes reward functions based on user query domains. Without fine-tuning, it improves accuracy on domain-specific evaluations.
With a pair of Mistral-7B models, Dr.SoW achieves a RewardBench score of 82.6, outperforming the best trained reward functions from same model class and demonstrating competitive performance against SoTA models in Safety (91.0) and Reasoning (88.0) domains. Further, we preference-tune Llama-3-8B-Instruct using data annotated by Dr.SoW. Our approach pushes Llama-3-8B to achieve a 37.4% ($+$15.1%) win rate on ArenaHard and a 40.7% ($+$17.8%) win rate on length-controlled AlpacaEval 2.0.
Machine Learning, ICML
Figure 1: We analyze how different model pairs $(\pi_{\text{strong}},\pi_{\text{weak}})$ impact the quality of the reward signal provided by (2). Each point represents one of 221 unique model pairs: 100 Llama-8B pairs (green) and 121 Mistral-7B pairs (blue). The x-axis denotes the alignment gap between $\pi_{\text{strong}}$ and $\pi_{\text{weak}}$, measured by ArenaHard scores, while the y-axis represents reward signal quality, measured by RewardBench scores. We observe a strong correlation between model alignment gap and reward signal quality, indicating that practitioners should pair a well-aligned $\pi_{\text{strong}}$ with a less-aligned $\pi_{\text{weak}}$ when using (2) as a reward signal.
@@ 1 Introduction
Preference tuning has advanced the capabilities of large language models (LLMs), but this progress relies on high-quality human preference data which is both costly and time-consuming to gather. Cutting-edge models are aligned with curated, quality-controlled human preference data, typically provided by specialized companies. While effective, this approach limits broader adoption due to prohibitive costs and limited transparency in data collection (Wang et al., 2024d). AI-feedback solutions are emerging as an alternative—either through a trained reward model (Dong et al., 2024) or proprietary LLM-as-a-judge (Cui et al., 2023). However, training such reward models still rely on costly initial human preference data; and LLM-as-a-judge approaches introduce licensing restrictions that generally prevent commercial use when using proprietary models.
We introduce Dr.SoW (Density Ratio of Strong-over-Weak), an automatic labeling method that not only drastically reduces manual costs in preference annotation, but also is comparable or beats proprietary model-as-a-judge method and trained reward models in reward accuracy and preference alignment outcome. Our method leverages the log-density ratio between a better-aligned and a less-aligned model to annotate preference data, offering a flexible approach applicable to any off-the-shelf open-source LLMs. Through extensive experiments across 221 model combinations (Figure 1), we provide design guidelines for selecting model pairs. Our findings reveal that a larger alignment gap between models enhances the reward signal for preference annotation, a principle we term the “Strong-over-Weak Hypothesis”. Our approach generalizes the DPO implicit reward, which restricts model pair selection to post-DPO and pre-DPO models (Chen et al., 2024). We demonstrate that by selecting a model pair with more significant alignment gap, the reward signal defined by Dr.SoW could outperform the DPO implicit reward (Figure 2). This flexibility allows models trained with diverse objectives—including SFT, RRHF, SLiC-HF, ORPO, SimPO, KTO, and IPO—to be used for data annotation. Moreover, our results offer actionable design guidelines for practitioners seeking to optimize reward function quality.
Customizing the reward function for data annotation is crucial to ensuring alignment with domain-specific needs. For instance, safety annotation may prioritize risk minimization and policy compliance, whereas code annotation might emphasize correctness and readability, and math annotation could focus on logical consistency and precision. A generic and one-size-fits-all reward function fails to capture these nuanced requirements. A common approach involves fine-tuning reward models for each domain, but this process is costly due to the need for domain-specific data collection and model training (Ji et al., 2024; Wang et al., 2024c). We streamline this process by introducing an end-to-end pipeline that identifies the domain of each user query and customizes the density-ratio reward function to prioritize relevant preference criteria. Specifically, Dr.SoW employs an adaptive router to classify queries into domains such as chat, reasoning, and safety. It then applies domain-specific instructions and in-context learning examples to refine preference criteria. In this way, we customize a density-ratio reward function from a general preference signal to domain-specific annotators. Experimental results show that adaptively customized density-ratio rewards significantly enhance both overall and domain-specific reward signal quality.
In summary, our main contributions are:
  • Cost-effective preference annotation. We introduce a scalable, cost-effective pipeline for preference data annotation. By leveraging the density ratio of off-the-shelf LLMs as a reward function, it drastically reduces the reliance on human annotation and allows for domain customization of reward without requiring additional data or fine-tuning. This automated annotation process can drastically lower the cost of human labeling, while also minimizing the expertise and computational resources traditionally needed for training reward models.
  • Broader model choice and better reward signals. Dr.SoW enables the use of any open-source or in-house models for preference data annotation. It goes beyond existing methods that rely on proprietary models or special model pairs for data annotation. We formalize the strong-over-weak hypothesis, which provides a principled guideline for selecting LLMs to produce a stronger reward signal. We observe that certain model pairs yield higher-quality reward functions than the DPO implicit reward.
  • Strong alignment performance. We provide an end-to-end preference data annotation pipeline and validate it through extensive experiments. With a pair of Mistral-7B models, Dr.SoW achieves a RewardBench score of 82.6, outperforming the best trained reward functions from same model class and demonstrating competitive performance against SoTA models in Safety (91.0) and Reasoning (88.0) domains. Further, we preference tune Llama-3-8B-Instruct using data annotated by Dr.SoW. Our approach pushes Llama-3-8B to achieve a 37.4% ($+$15.1%) win rate on ArenaHard and a 40.7% ($+$17.8%) win rate on length-controlled AlpacaEval 2.0. This outperforms model aligned with data from SoTA-level reward classifiers, proving our approach is both cost-effective and highly effective.
@@ 2 Background
Prior studies (Lin et al., 2024; Chen et al., 2024) has explored using implicit reward from direct policy optimization (DPO; Rafailov et al., 2023b) for preference data annotation. DPO is a preference-based fine-tuning method that does not require (explicit) reward modeling. Instead, it directly optimizes a policy language model $\pi_{\theta}$ using a reference model $\pi_{\text{ref}}$, typically an SFT model. The policy $\pi_{\theta}$ is initialized as $\pi_{\text{ref}}$, and the (implicit) reward function being optimized in DPO is:
$r_{\text{DPO}}(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}+\beta\log(Z(x))$ (1)
where $x$ is the prompt, $y$ is the answer, $\beta$ is a temperature hyperparameter and $Z(x)$ is a normalization constant. Ignoring the normalization constant, this reward function is the log-density ratio between a specific model pair: the policy model being optimized and its reference model.
A series of works (Lambert et al., 2024; Lin et al., 2024; Chen et al., 2024) explored leveraging the implicit reward function of DPO to annotate preference data. They proposed selecting a post-DPO model and a pre-DPO model to define a reward function. By definition, the pre-DPO model is the reference model (typically a SFT model) used during DPO training. Given a prompt $x$ and two responses, $y_{1}$ and $y_{2}$, the response with the higher reward is labeled as preferred, while the other is labeled as dispreferred.
@@ 3 Method
We study two research questions critical to density-ratio-based reward function design. First, we investigate whether alternative model pairs can produce stronger signals compared to the DPO implicit reward (section 3.1). Our experiments reveal a positive correlation between the alignment gap of model pairs (measured by the ArenaHard score) and the effectiveness of the reward function (evaluated through the RewardBench score). By increasing the gap in human alignment levels, we observe that certain model pairs yield a stronger reward signal than the DPO implicit reward. Second, we investigate whether we can further refine density-ratio reward based on domain characteristics of annotation data (section 3.2). We show that conditioning the density ratio with domain-related instructions and exemplars significantly improve overall and in-domain reward signal quality without requiring additional training.
@@
3.1 Density-ratio Reward Functions
@@
We explore constructing density-ratio-based reward function with various pairings of LLMs. At first glance, one might assume that the DPO model and its reference model would be the optimal pair for this purpose. To examine this hypothesis, we conduct an experiment using online iterative DPO (Xiong et al., 2023; Xu et al., 2023; Swamy et al., 2024) trained models from the Mistral and Llama-3 families. The key ideas of online iterative DPO training are: (1) the reference model is updated at each iteration (i.e., $\pi_{\text{ref}}=\pi_{\theta_{t-1}}$), and (2) the training data is also updated iteratively by sampling responses from $\pi_{\theta_{t-1}}(\cdot\mid x)$ and annotated with an external reward function.
In this online iterative DPO setting, the policy model $\pi_{\theta_{t}}$ at iteration $t$ uses the previous iteration’s policy model $\pi_{\theta_{t-1}}$ as its reference. According to the implicit DPO reward theory, one might expect the density ratio between $\pi_{\theta_{t}}$ and $\pi_{\theta_{t-1}}$ to provide an optimal reward function. However, Figure 2 shows that using weaker models—such as the base or SFT models—as the denominator in (2), instead of $\pi_{\theta_{t-1}}$, produces significantly better reward functions as evaluated by RewardBench. This finding indicates that the DPO implicit reward is empirically suboptimal compared with simply choosing weaker models in the denominator of (2), implication of which motivates us to propose the “Strong-over-Weak Hypothesis”.
@@
$r(x,y)=\log\frac{\pi_{\text{strong}}(y\mid x)}{\pi_{\text{weak}}(y\mid x)}.$ (2)
@@
Strong-over-Weak Hypothesis We conduct extensive experiments using $221$ distinct model pairs to construct various reward functions in (2) and evaluate their quality on RewardBench. Our findings reveal a strong correlation between the alignment gap of $\pi_{\text{strong}}$ and $\pi_{\text{weak}}$ and the effectiveness of the reward function, as quantified by the RewardBench score. As shown in Figure 1, achieving an effective reward function in (2) with a high RewardBench score requires a substantial human-alignment difference between $\pi_{\text{strong}}$ and $\pi_{\text{weak}}$. We refer to this insight as the “Strong-over-Weak Hypothesis”, which serves as a guiding principle for constructing density-ratio-based reward function as in (2). Our experiments span a range of models, including base, SFT, SimPO, KTO, ORPO, going beyond post-DPO and pre-DPO models (see Figure 4 for details). We summarize our key observations below.
@@
You are a helpful AI assistant. You follow the following guidelines when answering user questions. 1. Answer Constructive, Clear Questions
- Provide an answer when the user asks for factual information, constructive advice, or help with personal growth. Focus on offering practical, positive guidance.
2. Recognize Jokes, Puns, and Fictional Contexts
- Respond playfully when the question references humor, games, movies, or fictional scenarios. Acknowledge the fictional nature while keeping the tone light.
3. Avoid Answering Harmful, Illegal, or Malicious Questions
- Do not engage if the question promotes harm, illegal activities, or unethical behavior. Politely but firmly refuse to provide an answer, while keeping the response respectful.
4. Handle Sensitive Topics with Empathy
- Respond with care to questions about mental health, personal relationships, or emotionally charged situations. Acknowledge the user’s feelings, and offer general advice or suggest professional resources.